# The First and Second-Order Asymptotics of Covert Communication over AWGN Channels

### v4

@@ Abstract
This paper establishes the first- and second-order asymptotics of the maximal throughput under a covertness criterion defined by a Kullback-Leibler (KL) divergence bound $\delta$. Unlike prior works relying on discrete codebook designs, the achievability employs random coding with three key innovations: (1) dynamically scaled power constraints dependent on blocklength $n$, (2) codeword generation via truncated Gaussian distributions optimized for near-isotropy, and (3) codebook expurgation to ensure reliability. The analysis reveals three critical results. First, the maximal throughput scales as $O(n^{\frac{1}{2}})$ with second-order term $O(n^{\frac{1}{4}})$, where coefficients depend on the truncation hyper-parameter $\mu$ and the adversary’s noise variance $\sigma^{2}$. Second, the required secret key size exhibits a universal scaling law of $O(n^{\frac{1}{2}})$ even when the legitimate receiver’s channel quality marginally exceeds the adversary’s—a sharp contrast to discrete memoryless channels where such asymmetry reduces key requirements. Third, we prove tightness between achievability and converse bounds by optimizing $\mu$, demonstrating that truncated Gaussian codebooks achieve the information-theoretic limit. Theoretical contributions are further reinforced through an information-geometric perspective: truncated Gaussians minimize detectability by confining power within $O(n^{\frac{1}{2}})$ while maximizing throughput through near-capacity-achieving properties. The square-root power law is rigorously justified via asymptotic analyses of several divergences, unifying continuous and discrete signaling frameworks. These results advance covert communication design by providing explicit scaling laws, optimal codebook architectures, and a systematic methodology for balancing reliability, covertness, and resource constraints.
Index Terms:
Covert communication, channel coding rate, achievability and converse, KL divergence.
@@ I Introduction
This paper investigates the first- and second-order asymptotics of covert communication over AWGN channels, characterized as functions of blocklength $n$ decoding error probability $\epsilon$ and covertness parameter $\delta$, herein $\delta$ constrains the Kullback-Leibler (KL) divergence between the adversary’s observation distributions with and without communication. For achievability analysis, we extend our previous framework in [24] by maintaining the truncated Gaussian generating distribution with hyper-parameter $\mu$, ̵while replacing the sequential encoding scheme with an enhanced approach combining i.i.d. sampling and expurgation operations from [23]. Our reliability analysis introduces a novel decoding threshold selection mechanism that effectively bounds error probabilities for codewords with varying radii – ̵a distinctive feature not addressed in prior works [1, 2]. The channel resolvability analysis advances in two key aspects: First, ̵we circumvent the quasi-metric limitations of KL divergence in Euclidean space by directly analyzing the divergence between the Gaussian mixture induced by codewords and the noise distribution, avoiding reliance on modified triangle inequalities applicable to DMCs. Second, ̵we establish an explicit connection between adversary noise levels and required key bits, revealing a fundamental dichotomy between AWGN and DMC scenarios. ̵Notably, the key size remains $O(\sqrt{n})$ even when adversary noise marginally exceeds main channel noise – a critical divergence from DMC behavior that we rigorously explain through channel characteristic analysis. We further derive tight bounds for system performance: The supremum of achievable throughput and infimum of key bits under optimal $\mu$ selection precisely match the converse bounds, demonstrating theoretical completeness. ̵The optimality of truncated Gaussian distributions is justified through local information geometry principles [25, 26], where covertness realization corresponds to positioning the signal distribution within a quasi-$\varepsilon$ neighborhood of noise distribution in statistical manifold space. ̵This geometric interpretation bridges coding theory with machine learning applications [27, 28]. The square root law governing power scaling $\Psi(n)=O(\frac{1}{\sqrt{n}})$ emerges naturally from asymptotic analyses of Gaussian divergence measures and noise distribution approximation accuracy.
@@
  • Complete asymptotic characterization: First- and second-order expansions for Gaussian random coding in AWGN covert communication, explicitly parameterized by $(n,\epsilon,\delta,\mu)$. ̵The $\mu$-dependence underscores finite-blocklength power constraint necessities.
@@
  • Converse bound analysis: Matching achievability supremum and converse bounds under optimal $\mu$ selection, establishing fundamental performance limits.
@@
The rest of this paper is organized as follows. In Section II, we present some preliminaries on notations, the channel model, the definitions of codes and some distributions. Section III provides some technical preparations, mainly the properties of related distributions. Section IV presents the first and second order asymptotics of the achievability bound. Section V presents the investigation of the converse bound. The framework of local information geometry to explain the optimality of truncated Gaussian distribution is provided in Section VI. The investigation on the connection between Square Root Law, the choice of the power level and the asymptotic behavior of several divergences of Gaussian distributions is provided in Section VII. Section VIII concludes the paper.
@@ II-A Notations
[…]
@@ II-B The Channel Model
The communication model considered in this work is illustrated in Fig.1, where a transmitter Alice communicates with a legitimate receiver Bob, and the signals are also observed by an adversary Willie. The channel between Alice and Bob, Willie are $W_{1}$ and $W_{2}$, which are two AWGN channels modeled as
@@
The messages are indexed by $w\in\langle M\rangle$. A codeword over the AWGN channel is represented by a vector in $\mathbb{R}^{n}$: $\bm{x}=(x_{1},x_{2},\cdots,x_{n})$. With channel transition models $W_{1}(y|x)$ and $W_{2}(z|x)$, we have $W^{n}_{1}(y^{n}|x^{n})=\prod_{i=1}^{n}W_{1}(y_{i}|x_{i})$ and $W^{n}_{2}(z^{n}|x^{n})=\prod_{i=1}^{n}W_{2}(z_{i}|x_{i})$, which can also be expressed as $\bm{y}=\bm{x}+\bm{n}_{0}$ and $\bm{z}=\bm{x}+\bm{n}_{1}$. The communication may be assisted by a uniformly distributed shared secret key $S\in\langle K\rangle$, which is kept secret from the adversary. Thus, the encoder and decoder are represented by
$f:\langle K\rangle\times\langle M\rangle\rightarrow\mathbb{R}^{n}\,\ \,\text{and}\,\ \,\phi:\mathbb{R}^{n}\times\langle K\rangle\rightarrow\langle M\rangle.$ (7)
The transmission between Alice and Bob results in a mixture Gaussian distribution at the adversary induced by the codewords $\bm{x}_{sw},s=1,\cdots,K$ and $w=1,\cdots,M$ through the channel $W_{2}$:
$\hat{Q}^{(n)}_{\bm{Z}}(\bm{z})=\sum_{s=1}^{K}\sum_{k=1}^{M}\frac{1}{MK}W^{n}_{2}(\bm{z}|\bm{x}_{sw})$ (8)
where $\bm{x}_{sw}=f(s,w)$ is a codeword. When there is no transmission, the distribution of channel noise at the adversary with blocklength $n$: $\hat{Q}^{(n)}_{0}(\bm{z})\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I}_{n})$, the probability density function is denoted as $\hat{f}_{0}^{(n)}(\bm{z})$.
There are two requirements in the transmitter’s objective:
  1. 1.
    She should communicate a uniformly distributed message $w\in\langle M\rangle$ reliably, i.e., the legitimate receiver should has a small probability $\epsilon$ of decoding error.
  2. 2.
    The adversary should has low probability on successful detection of the communication, and the detection is performed by binary hypothesis testing between two distributions $\hat{Q}^{(n)}_{\bm{Z}}$ and $\hat{Q}^{(n)}_{0}$.
The detection performance of the adversary is usually measured by $\mathbb{V}(\hat{Q}^{(n)}_{\bm{Z}},\hat{Q}^{(n)}_{0})$ [29]. When it is close to 00, it is generally believed that no detector at Willie can effectively discriminate the induced output distribution and the distribution of noise, thus cannot distinguish whether Alice is communicating with Bob. Nevertheless, it is more convenient to use KL divergence to measure the discrimination between $\hat{Q}^{(n)}_{\bm{Z}}$ and $\hat{Q}^{(n)}_{0}$ in the asymptotic setting as a metric of covertness. In this work, we follow the conventions in the works [13][14] to use the following covert constraint:
$\mathbb{D}(\hat{Q}^{(n)}_{\bm{Z}}\|\hat{Q}^{(n)}_{0})\leq\delta.$ (9)
@@ II-C Definitions for Codes over AWGN Covert Channels
A $(n,M,\epsilon,\Psi)_{m}$ code $\mathcal{C}=\{c_{1},c_{2},\cdots,c_{M}\}$ for a AWGN channel consists of a message set $W\in\mathcal{W}=\{1,...,M\}$, an encoder at the transmitter Alice $f_{n}:\mathcal{W}\rightarrow\mathcal{C},w\mapsto\bm{x}$, and a decoder at the legitimate user Bob $g_{n}:\mathbb{R}^{n}\rightarrow\mathcal{W},\bm{y}\rightarrow\hat{w}$ with the condition that the encoder and decoder pair satisfies
@@
    maximal power constraint: each codeword $c_{i}$ satisfies $\|c_{i}\|^{2}\leq n\Psi$.
If the maximal power constraint is substituted by average power constraint: the codebook satisfies $\frac{1}{M}\sum_{i=1}^{M}\|c_{i}\|^{2}\leq n\Psi$, then it is a $(n,M,\epsilon,\Psi)_{a}$ code.
Fig. 1: The channel model of Gaussian LPD communication in Section II
As the adversary should have low probability of detection on the communication, it puts additional constraint on the codebook. The following definitions are for codes over AWGN covert channels.
@@
  • A $(n,M,\epsilon,\Psi_{max})_{m}$ code is a $(n,M,\epsilon,\Psi_{max},\delta)_{m}$ code if it also satisfies (9);
  • A $(n,M,\epsilon,\Psi_{avg})_{a}$ code is a $(n,M,\epsilon,\Psi_{avg},\delta)_{a}$ code if it also satisfies (9).
The maximal sizes of $(n,M,\epsilon,\Psi_{max})_{m}$, $(n,M,\epsilon,\Psi_{avg})_{a}$, $(n,M,\epsilon,\Psi_{max},\delta)_{m}$, $(n,M,\epsilon,\Psi_{avg},\delta)_{a}$ codes are denoted by $M^{*}_{m}(n,\epsilon,\Psi_{max})$, $M^{*}_{a}(n,\epsilon,\Psi_{avg})$, $M^{*}_{m}(n,\epsilon,\Psi_{max},\delta)$ and $M^{*}_{a}(n,\epsilon,\Psi_{avg},\delta)$, respectively.
If $\Psi_{max}=\Psi_{avg}=\Psi_{lim}$, we have the following inequalities:
$M^{*}_{m}(n,\epsilon,\Psi_{lim},\delta)\leq M^{*}_{m}(n,\epsilon,\Psi_{lim})\leq M^{*}_{a}(n,\epsilon,P_{lim}),$ (11)
$M^{*}_{m}(n,\epsilon,\Psi_{lim},\delta)\leq M^{*}_{a}(n,\epsilon,\Psi_{lim},\delta)\leq M^{*}_{a}(n,\epsilon,\Psi_{lim}).$ (12)
Note that different from DMCs, there is always power constraints for codebooks over AWGN channels. In general, the covert constraint will give rise to another average power constraint $\Psi(n)$ since the covert constraint is measured on the signals of blocklength $n$, which may make original power constraint $\Psi_{max}$ or $\Psi_{avg}$ redundant. For convenience, it is assumed that the natural power constraint of the channel is redundant, i.e., the channel itself is not an obstruction for Alice to seek a code to satisfy the covert constraint (9). In addition, following our previous works [3][24], the maximal probability of error is considered as the reliability metric. Thus, the throughput of covert AWGN channels has the following definitions.
@@
    Achievability: A lower bound on the maximal throughput $M^{*}(n,\epsilon,\delta)$ satisfying covert constraint (9) that can be guaranteed to exist with given blocklength $n$, and maximal probability of error $\epsilon$.
@@
    Converse: An upper bound on the maximal throughput $M^{*}(n,\epsilon,\delta)$ satisfying covert constraint (9) with given blocklength $n$, and maximal probability of error $\epsilon$.
II-D Some Distributions

### v5

@@ Abstract
This paper establishes the first- and second-order asymptotics of the maximal throughput under a covertness criterion defined by a Kullback-Leibler (KL) divergence bound $\delta$. The achievability employs random coding with three key innovations: (1) dynamically scaled power constraints dependent on blocklength $n$, (2) codeword generation via truncated Gaussian distributions optimized for near-isotropy, and (3) codebook expurgation to ensure reliability. The analysis reveals two critical results. First, the maximal throughput scales as $O(n^{\frac{1}{2}})$ with second-order term $O(n^{\frac{1}{4}})$. Compared with the previous results, we present analytic formulas of the asymptotics as functions of the truncation hyper-parameter $\mu$, the adversary’s noise variance $\sigma^{2}$ and the upper bound $\delta$ on KL divergence. Second, the required secret key size exhibits a universal scaling law of $O(n^{\frac{1}{2}})$ even when the legitimate receiver’s channel quality marginally exceeds the adversary’s—a sharp contrast to discrete memoryless channels where such asymmetry reduces key requirements. Theoretical contributions are further reinforced through an information-geometric perspective: truncated Gaussians minimize detectability by confining power within $O(n^{\frac{1}{2}})$ while maximizing throughput through near-capacity-achieving properties. The square-root power law is further justified via asymptotic analyses of several divergences of Gaussian family with different variances. These results advance covert communication design by providing explicit scaling laws, optimal codebook architectures, and a systematic methodology for balancing reliability, covertness, and resource constraints.
@@ I Introduction
This paper investigates the first- and second-order asymptotics of covert communication over AWGN channels, characterized as functions of blocklength $n$ decoding error probability $\epsilon$, the noise level at the adversary $\sigma$ and covertness parameter $\delta$, herein $\delta$ constrains the Kullback-Leibler (KL) divergence between the adversary’s observation distributions with and without communication. For achievability analysis, we extend our previous framework in [23] by maintaining the truncated Gaussian generating distribution with hyper-parameter $\mu$, ̵while replacing the sequential encoding scheme with an enhanced approach combining i.i.d. sampling and expurgation operations from [22]. Our reliability analysis introduces a novel decoding threshold selection mechanism that effectively bounds error probabilities for codewords with varying radii – ̵a distinctive feature not addressed in prior works [20, 24]. The channel resolvability analysis advances in two key aspects: First, ̵we circumvent the quasi-metric limitations of KL divergence in Euclidean space by directly analyzing the divergence between the Gaussian mixture induced by codewords and the noise distribution, avoiding reliance on modified triangle inequalities applicable to DMCs. Second, ̵we establish an explicit connection between adversary noise levels and required key bits, revealing a fundamental dichotomy between AWGN and DMC scenarios. ̵Notably, the key size remains $O(\sqrt{n})$ even when adversary noise marginally exceeds main channel noise – a critical divergence from DMC behavior that we rigorously explain through channel characteristic analysis. The optimality of truncated Gaussian distributions is justified through local information geometry principles [26, 27], where covertness realization corresponds to positioning the signal distribution within a quasi-$\varepsilon$ neighborhood of noise distribution in statistical manifold space. ̵This geometric interpretation bridges coding theory with machine learning applications [28, 29]. The square root law governing power scaling $\Psi(n)=O(\frac{1}{\sqrt{n}})$ emerges naturally from asymptotic analyses of Gaussian divergence measures and noise distribution approximation accuracy. Note that our work only considers AWGN channel model and do not take into account of the signal duration or spectral width. Under such circumstance, the transmission involves specific modulation, and has been investigated in [30][31].
@@
  • Complete asymptotic characterization: First- and second-order expansions for Gaussian random coding in AWGN covert communication, explicitly parameterized by $(n,\epsilon,\sigma,\delta,\mu)$. ̵The $\mu$-dependence underscores finite-blocklength power constraint necessities.
@@
The rest of this paper is organized as follows. In Section II, we present some preliminaries on notations, the channel model, the outline of the proof of the achievability and the converse, the definitions of codes and some distributions. Section III presents the main results, which include the first and second order asymptotics of the achievability bound and the investigation of the converse bound. The framework of local information geometry to explain the optimality of truncated Gaussian distribution is provided in Section IV. The investigation on the connection between Square Root Law, the choice of the power level and the asymptotic behavior of several divergences of Gaussian distributions is provided in Section V. Section VI concludes the paper.
@@ II-A Notations
Throughout the paper, $\log$ denotes logarithmic function with base $2$. We use upper case letters, such as $X$, for random variables and lower case letters, such as $x$ for their realizations. Moreover, upper case letters with boldface fonts are used for random vectors, e.g., $\bm{Z}$, and their realizations are denoted with lower case letters of boldface font, e.g., $\bm{z}$. The norm $\|\cdot\|$ will be used for $l^{2}$ norm $\|\cdot\|_{2}$. We use * for convolution operation and $\langle M\rangle$ for the set $\{1,\cdots,M\}$. If $a$ is a real number, then $|a|$ denotes its absolute value. If $\mathcal{X}$ is a set, then $|\mathcal{X}|$ denotes its cardinality. The function $\Phi$ denotes the cumulative distribution function of standard normal random variable. For two distributions $P$ and $Q$ in the same probability space, some related divergences in this paper are defined as follows.
@@ II-B The Channel Model
The AWGN covert channel model considered in this work is illustrated in Fig.1. A transmitter Alice communicates with a legitimate receiver Bob through a main channel $W_{1}$, and the signals are also observed by an adversary Willie through another channel $W_{2}$, which are modeled as
@@
The messages are indexed by $w\in\langle M\rangle$. A codeword over the AWGN channel is represented by a vector in $\mathbb{R}^{n}$: $\bm{x}=(x_{1},x_{2},\cdots,x_{n})$. With channel transition models $W_{1}(y|x)$ and $W_{2}(z|x)$, we have $W^{n}_{1}(y^{n}|x^{n})=\prod_{i=1}^{n}W_{1}(y_{i}|x_{i})$ and $W^{n}_{2}(z^{n}|x^{n})=\prod_{i=1}^{n}W_{2}(z_{i}|x_{i})$.
@@ II-C Definitions for Codes over AWGN Covert Channels
A $(M,n,\epsilon,\Psi)_{m}$ code $\mathcal{C}=\{\bm{c}_{1},\bm{c}_{2},\cdots,\bm{c}_{M}\}$ for a AWGN channel (5), consists of a message set $W\in\langle M\rangle$, an encoder at the transmitter Alice $f_{n}:\mathcal{W}\rightarrow\mathbb{R}^{n},w\mapsto\bm{x}$, and a decoder at the legitimate user Bob $g_{n}:\mathbb{R}^{n}\rightarrow\langle M\rangle,\bm{y}\rightarrow\hat{w}$ with the condition that the encoder and decoder pair satisfies
@@
    maximal power constraint: each codeword $\bm{c}_{i}$ satisfies
    $\|\bm{c}_{i}\|^{2}\leq n\Psi,$ (8)
If the maximal power constraint is substituted by average power constraint: the codebook satisfies
$\frac{1}{M}\sum_{i=1}^{M}\|\bm{c}_{i}\|^{2}\leq n\Psi,$ (9)
then it is a $(M,n,\epsilon,\Psi)_{a}$ code.
Fig. 1: The model of covert communication over AWGN channels in Section II
For AWGN covert channels, the following formal definition characterizes the codes applicable for them.
@@
An $(M,K,n,\epsilon,\sigma,\delta)$ code for AWGN covert channels (5) and (6) consists of an encoder/decoder pair
$f:\langle K\rangle\times\langle M\rangle\rightarrow\mathbb{R}^{n}\,\ \ \,\text{and}\,\ \ \ \ \,\phi:\mathbb{R}^{n}\times\langle K\rangle\rightarrow\langle M\rangle,$
with
$\bm{x}_{sw}=f(s,w)\,\ \ \,\text{and}\,\ \ \ \ \,\hat{w}=\phi(\bm{y},s)$ (10)
The codewords $\bm{x}_{sw},s\in\langle K\rangle,w\in\langle M\rangle$ compose a codebook satisfies the following properties.
@@
    Each message $w\in\langle M\rangle$ should be transmitted reliably:
    $P_{err}\triangleq\underset{s\in\langle K\rangle,w\in\langle M\rangle}{\max}\mathbb{P}(\phi(\bm{Y},s)\neq W|S=s,W=w)\leq\epsilon,$ (11)
@@
    The adversary should have low probability on successful detection of the communication, which leads to the KL divergence requirement on the distribution $\hat{Q}^{(n)}_{\bm{Z}}$ with respect to $\hat{Q}^{(n)}_{0}$:
    $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})\leq\delta,$ (12)
    where $Q^{(n)}_{\bm{Z},\bm{C}}$ is a mixture Gaussian distribution induced by the codebook at the adversary:
    $Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})=\sum_{s=1}^{K}\sum_{w=1}^{M}\frac{1}{MK}W^{n}_{2}(\bm{z}|\bm{x}_{sw}).$ (13)
    and $Q^{(n)}_{\bm{Z},0}$ is the distribution of channel noise at the adversary with blocklength $n$: $Q^{(n)}_{\bm{Z},0}(\bm{z})\sim\mathcal{N}(\bm{0},\sigma^{2}\bm{I}_{n})$.
Note that from Definition 1, there is always power constraint for codebooks over AWGN channels. For codes for the AWGN covert channel, there are two additional ingredients. (1), The noise level $\sigma$ of the channel $W_{2}$ plays an important role and will affect the power level of legitimate codes because the covert constraint (12) depends on the parameter $\sigma$; (2), To accomplish the target of covertness, some shared secret key $S$ may be necessary. To take the power level of codes into consideration, we introduce the following extension for AWGN covert codes by combining Definition 1 and Definition 2.
Definition 3.
  • An $(M,K,n,\epsilon,\Psi,\sigma,\delta)_{m}$ code is a $(M,K,n,\epsilon,\Psi,\sigma,\delta)_{m}$ code if for each $s\in\langle K\rangle$, the sub-code $\{\bm{x}_{sw},w\in\langle M\rangle\}$ satisfies maximal power constraint:
    $\|\bm{x}_{sw}\|^{2}\leq n\Psi\,\ \,\text{holds for each}\,\ \,w\in\langle M\rangle$ (14)
  • An $(M,K,n,\epsilon,\Psi,\sigma)_{a}$ code is a $(M,K,n,\epsilon,\Psi,\sigma,\delta)_{a}$ code if for each $s\in\langle K\rangle$, the sub-code $\{\bm{x}_{sw},w\in\langle M\rangle\}$ satisfies average power constraint:
    $\frac{1}{M}\sum_{w=1}^{M}\|\bm{x}_{sw}\|^{2}\leq n\Psi\,\ \,\text{holds for each}\,\ \,w\in\langle M\rangle$ (15)
The maximal sizes of $(M,n,\epsilon,\Psi)_{m}$, $(M,n,\epsilon,\Psi)_{a}$, $(M,K,n,\epsilon,\Psi,\sigma,\delta)_{m}$, $(M,K,n,\epsilon,\Psi,\sigma,\delta)_{a}$, $(M,K,n,\epsilon,\sigma,\delta)$ codes are denoted by $M^{*}_{m}(n,\epsilon,\Psi)$, $M^{*}_{a}(n,\epsilon,\Psi)$, $M^{*}_{m}(n,\epsilon,\Psi,\sigma,\delta)$, $M^{*}_{a}(n,\epsilon,\Psi,\sigma,\delta)$ and $M^{*}(n,\epsilon,\sigma,\delta)$ respectively. In addition, for any $(M,K,n,\epsilon,\sigma,\delta)$ code, the corresponding number of key bits is denoted as $K(n,\epsilon,\sigma,\delta)$.
Remark 1.
In this work, we follow our previous works [7][23] and [22], and choose the maximal probability of error as the reliability metric.
Remark 2.
The detection performance of the adversary is usually measured by $\mathbb{V}(Q^{(n)}_{\bm{Z},\bm{C}},Q^{(n)}_{\bm{Z},0})$ [34]. When it is close to 0, it is generally believed that no detector at Willie can effectively discriminate the induced output distribution and the distribution of noise, thus cannot distinguish whether Alice is communicating with Bob. Nevertheless, it is more convenient to use KL divergence to measure the discrimination between $Q^{(n)}_{\bm{Z},\bm{C}}$ and $Q^{(n)}_{\bm{Z},0}$ in the asymptotic setting as a metric of covertness. In this work, we follow the conventions in the works [11][12] to use (12) as covert constraint.
Remark 3.
As the distribution $Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})$ depends on the codewords $\bm{x}_{sw}$ and the codewords are randomly generated, the quantity $Q^{(n)}_{\bm{Z},\bm{C}}(\bm{z})$ itself is a random variable.
For the above quantities, we have the following lemma, the proof of which can be found in Appendix A.
Lemma 1.
$M^{*}_{m}(n,\epsilon,\Psi,\sigma,\delta)\leq M^{*}_{m}(n,\epsilon,\Psi)\leq M^{*}_{a}(n,\epsilon,\Psi),$ (16)
$M^{*}_{m}(n,\epsilon,\Psi,\sigma,\delta)\leq M^{*}(n,\epsilon,\sigma,\delta),$ (17)
$M^{*}_{m}(n,\epsilon,\Psi,\sigma,\delta)\leq M^{*}_{a}(n,\epsilon,\Psi,\sigma,\delta)\leq M^{*}(n,\epsilon,\sigma,\delta).$ (18)
The aim of this work is to characterize the quantity $M^{*}(n,\epsilon,\sigma,\delta)$ through the following two parts.
  1. 1.
    Achievability: A lower bound of $M^{*}(n,\epsilon,\sigma,\delta)$ is obtained by constructing a code satisfying (11) and (12) by some coding scheme, usually through random coding argument. Meanwhile, the corresponding number of key bits is denoted as $K(n,\epsilon,\sigma,\delta)$.
  2. 2.
    Converse: an upper bound of $M^{*}(n,\epsilon,\sigma,\delta)$ is provided such that any code satisfying (11) and (12) should have size larger than it.
II-D Outline of The Proof
In this section, we provide a road map on the proof of the achievability and converse bounds.
There are mainly four steps for achievability.
  1. 1.
    Random coding scheme, decoding method, which are introduced in Section III-A. A code family consists of $MK$ codewords $\bm{x}_{sw},s=1,\cdots,K,w=1,\cdots,M$ are randomly generated according to truncated Gaussian distribution, and threshold decoding is adopted.
  2. 2.
    The reliability analysis (also in Section III-A) focuses on bounding two types of decoding error probability $\varepsilon_{sw}^{(1)}$ and $\varepsilon_{sw}^{(1)}$. The maximum of the first type is bounded in Lemma 4. The maximum of the second type can not be directly bounded. The expectation of it is bounded in Lemma 6, which leaves the bounding the maximum of the latter type for follow-up analysis.
  3. 3.
    The resolvability analysis in Section III-B focuses on bounding $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$. It also can not be directly bounded, and we provide an upper bound of its expectation in Theorem 4. Lemma 7, Lemma 8 and Lemma 9 are necessary preliminary results for Theorem 4.
  4. 4.
    The existence and characterization of the code family are provided in Section III-C. The expurgation operation is adopted to bound the maximum of the second type of error for each sub-code, meanwhile, the quantity $\mathbb{D}(Q^{(n)}_{\bm{Z},\bm{C}}\|Q^{(n)}_{\bm{Z},0})$ for the whole remaining codewords is shown to be low. The proof relies on the application of McDiarmid’s Inequality on codewords over AWGN channels.
The proof of the converse consists of two steps.
  1. 1.
    We show the converse constraint leads to an upper bound for the average power of the code family in Lemma 13.
  2. 2.
    By using the pigeon-hole principle, the converse bound for each sub-code under maximal power constraint is derived in Corollary 3, which further leads to characterization of the converse bound under covert constraint in Theorem 7.
II-E Related Distributions